OpenStudy (anonymous):

what is the trig solution to 2cosxsinx-cosx=0

OpenStudy (anonymous):

you have 2 ways 1. \[\cos x(2\sinx-1)=0\] 2. \[\sin {2x}=\sin(\frac{\pi}{2}-x)\]

OpenStudy (anonymous):

@Disco94 , you got it?

OpenStudy (anonymous):

what???

OpenStudy (anonymous):

Sorry. :)

OpenStudy (mertsj):

Factor out cos x \[\cos x(2\sin x-1)=0\] \[\cos x=0 \] or \[\sin x=\frac{1}{2}\]

OpenStudy (anonymous):

\[ 2 \cos x \sin x- \cos x=0 \implies \cos x (2 \sin x -1) = 0 \] \[\cos x = 0 \implies x = (2n+1) \frac \pi 2, n \in \mathbb{W} \] or \[ \sin x = \frac 12 \implies x = n\pi +(-1)^n \frac \pi 6, n \in \mathbb{W} \]

OpenStudy (anonymous):

the answer is the cosx= pie/2 and 3pie/2 the sinx=1/2 and pie/6 and 5pie/6 but i need to know how to get the answer

OpenStudy (mertsj):

So the simple answer is \[x=\frac{pi}{2}, \frac{3pi}{2}, \frac{pi}{6}, \frac{5pi}{6}+2n pi\]

OpenStudy (anonymous):

Put \( n = 0,\pm 1 \) in the equation in my derivation you will get your desired answer. But I have to say the question is not clear.